Goddard Company has used the FIFO method of inventory valuation since it began operations in 2018. Goddard decided to change to the average cost method for determining inventory costs at the beginning of 2021. The following schedule shows year-end inventory balances under the FIFO and average cost methods:

Year FIFO Average Cost
2018 $45,700 $55,400
2019 $80,100 $71,700
2020 $85,800 $80,100

Required:
1. Ignoring income taxes, prepare the 2021 journal entry to adjust the accounts to reflect the average cost method.
2. How much higher or lower would cost of goods sold be in the 2020 revised income statement?
Answer #1

1)

Retained earnings 5,700
Inventory 5,700

(85,800 - 80,100)

2)

Decrease in beginning inventory = (80,100-71,700) = 8,400

Decrease in ending inventory = (85,800-80,100) = 5,700

Cost of Goods Sold will be lower by 2,700